what if the classes bought 120 containers and had 30 more cans than bags how many bags and how many cans and bags did they buy?
Answer by LinnW(1048)

set c = number of cans
b = number of bags
c + b = 120
c = b + 30 we need to add 30 to number of bags to make these equal
If we substitute ( b+30 ) for c in
c + b = 120 we have
(b+30) + b = 120
2b + 30 = 120
Subtract 30 from each side
2b = 120 - 30
2b = 90
b = 45
Since c = b+20
c = 45+20
c = 65